A 60.0L tank of chlorine gas at 27°C and 125atm springs a leak. When the leak was discovered, the pressure was reduced to 50atm. How many moles of chlorine gas escaped?
1
Expert's answer
2020-10-19T14:02:18-0400

Let's solve the Clapeyron-Mendeleev equation by equation:

"P1V=n1RT"

"P2V=n2RT"

"(P1-P2)V=(n1-n2)RT"

"\Delta PV=\Delta nRT"

"\Delta n=\frac{\Delta PV}{RT}"

"\Delta n=\frac{\Delta PV}{RT}=\frac{(125-50)60}{0.082*300}=\frac{4500}{24.6}=182.93" mol